State whether the following statement true or false

QUESTION 4

The number of arrivals per minute at a bank located in the central business district of a large city

was recorded over a period of 200 minutes, with the following results:

Arrivals Frequency

0 14

1 31

2 47

3 41

4 29

5 21

6 10

7 5

8 2

The probability of at least two arrivals per minute at the bank is 0:155:


1
Expert's answer
2022-05-20T12:09:54-0400
"P(X\ge2)=1-P(X=0)-P(X=1)"

"=1-\dfrac{14+31}{200}=0.775\not=0.155"